Jen Butler has been pricing Speed-Pass train fares for a group trip to New York Three adults and four children must pay $120. Two adults and three children must pay $85. Find the price of the adults ticket and the price of a child's ticket.
Answer by jorel1380(3719)

Let a be adults, and c be children. Then:
3a+4c=120
2a+3c=85
So:
6a+8c=240
6a+9c=255
c=15
a=20
Adults are $20 each, and children are $15 each. ☺☺☺☺